The curves result from structure function parametrizations which cover the transition through the hadronic mass scale region in photon virtuality [?]. Further information on these cross sections and their interpretation can obtained from the talk by A. Bornheim. Here we point out first the interesting threshold behavior for x > 0.06, where the proton diameter exceeds the virtual photon lifetime in the the proton rest frame. The origin of the steep energy dependence at high Q^{2} and high W, the HERA discovery, is clearly in need of an independent explanation. The cross section rises much more steeply than that for quasireal photons, where the energy dependence is comparable to that of the total hadronic cross sections. The behavior of these cross sections at energies higher than those reached at HERA remains an open and intriguing question.^{1}